OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (anonymous):

@Hero @perl

OpenStudy (anonymous):

@Vitas

OpenStudy (anonymous):

@Nairz

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

@Vitas Can you help me please

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

please help @Haseeb96

OpenStudy (anonymous):

4x-3y=-12 Organizing 4x+12=3y (4x+12)/3=y Slope is 4/3 Writing the line y=4x/3+n -3,4 = x,y 4=4/3*-3+n n=8 and line becomes y=4x/3+8

OpenStudy (anonymous):

Thanks i thought you were not there Lol @Vitas

OpenStudy (anonymous):

Can you help with another @Vitas

OpenStudy (anonymous):

Your welcome :) of course I can

OpenStudy (anonymous):

ok ill open another